OpenStudy (anonymous):

If T: R3 -> R3 is a linear transformation such that T1 [1,0,0] = [-3,2,-2], T2 [0,1,0] = [4,-2,2], and T3 [0,0,1] = [-4,3,1], the T [0,-3,4] =

OpenStudy (zzr0ck3r):

\(T(\begin{bmatrix} 0 \\ -3 \\ 4 \end{bmatrix}) = \begin{bmatrix} -3 & 4 & -4 \\ 2 & -2 & 3 \\ -2 & 2 & 1 \end{bmatrix} *\begin{bmatrix} 0 \\ -3 \\ 4 \end{bmatrix}= \ \huge?\)

OpenStudy (zzr0ck3r):

@atercero

OpenStudy (experimentx):

this shouldn't be difficult ... they are [1,0,0], [0,1,0], [0,0,1] spans R^3 ... find a, b ,c ... such that a[1,0,0]+b[0,1,0]+c [0,0,1] = [0,-3,4] then use the definition of Linearity.